# KNARsack: Teaching Neural Algorithmic Reasoners to Solve Pseudo-Polynomial Problems

### v1

@@ 1 Introduction
Knapsack is NP-hard. [4] CLRS-30 omits this class of problems, due to the impossibility of generating sufficiently many samples for them (unless NP=co-NP). [5] However, for Knapsack, there exists a dynamic programming solution whose runtime depends on the numeric value of the capacity rather than its bit-length. This makes Knapsack a pseudo-polynomial problem [^1] Informal term meaning the problem admits a pseudo-polynomial algorithm; formally, it is weakly NP-hard., lifting the implications of Yehuda et al. [5], provided we do not consider extreme capacities. Despite its theoretical and practical significance, the problem has not been explored within the NAR framework. Thus, we develop a neural algorithmic reasoner for Knapsack following a two-phase pipeline: dynamic programming table construction and solution reconstruction, as shown in Figure 1. Such approach is transferable to other pseudo-polynomial problems, as further discussed in Appendix B.
@@ 2.2 Pseudo-Polynomial Modeling
where $\delta$ is the Kronecker delta function and $M=10$ is the cutoff. Such categorical encoding gives the model the ability to choose the correct past states that influence the current one. Figure 2 shows the impact of EL on an in-distribution problem instance in comparison to standard random node positional encoding. For more details, see Appendix E.
We additionally observe that the CLRS-30’s processors struggle to handle the Knapsack DP scalar values larger than those seen in training. We propose to mitigate this issue by adopting a homogenous processor [8], which enforces the model to be invariant to the scale of the item values. This substantially improves the value generalization performance of the model on instances larger than those seen in training. Appendix F provides more results and a discussion.
@@
where $\bm{p}$} and $\bm{w}$} are the vectors of positional encodings and weights (capacity nodes are padded with 0), $0\leq c,c^{\prime}\leq C$ and $C+1\leq i\leq C+n$ and ${\mathbb{I}}_{\mathcal{G}}=\emptyset$. To model the reconstruction process, the hints provide information on the item currently under consideration, the total remaining capacity, and the set of items selected up to that point. For each item, the model outputs the probability that it belongs to the solution. Due to the space constraints, details can be found in Appendix D.
To reconstruct the solution from the $\mathrm{decision}$ table, it is sufficient to have information about the item weights. During reconstruction, as shown in Figure 3, having item values leads to worse generalization, as the model is trying to find a shortcut to the solution, without actually learning to navigate the $\mathrm{decision}$ table. We hypothesise that this is the reason why the joint model approach did not work. We also note that having two alternating steps, one for decision and one for moving the pointer works better than combining the two. Further details are provided in Figure 7, Appendix D.
@@
An alternative would be to perform reconstruction based on the $\mathrm{decision}$ probability table, as it allows for end-to-end differentiability. In Appendix G, we describe in more detail such a differentiable deterministic reconstruction. Sadly, our initial experiments in training an NAR construction, coupled with deterministic reconstruction, were unstable. Nevertheless, in the evaluation, we include as a baseline a model that, during the inference phase, performs deterministic reconstruction based on the $\mathrm{decision}$ table generated by NAR construction model.
@@ 3 Evaluation
Our experiments are implemented on top of CLRS-30 [7] (3 seeds for std). Hyperparameters (learning rate, optimiser, etc.) are inherited, including the use of Triplet-GMPNN processor [9] as well as the standard node positional encoding. Differently from CLRS-30, we use the model from the final epoch for testing, as it was recommended in [10]. Furthermore, during the construction training phase, we perform 30,000 epochs, as we observed that the model does not stabilize within the first 10,000 epochs. The model is trained on samples where $n\leq 16$ and $C\leq 16$ (in-distribution). The out-of-distribution we consider different values of $n$ and $C$ (up to $n=64$ and $C=64$). Each configuration consist of 64 samples. Item weights are uniformly sampled from $\{1,2,\ldots,8\}$, and item values are uniformly sampled from the interval $[0,1]$. Following CLRS-30, during testing, we report the micro-F1 score. We also add exact-match accuracy – the fraction of samples whose discretized solution, obtained by greedily selecting the highest-probability item until capacity is reached, exactly matches the ground truth.
@@
Alongside the models with regular processors (reg. c. + reg. r.), and a homogenous processor during construction (homo. c. + reg. r.), we evaluate a no-hint baseline, which directly predicts the optimal subset from the input. It is implemented using the no-hint mode in CLRS-30 with $2n$ message-passing steps, and includes scalar features for item weights and capacity. We additionally zero-shot evaluate the homogenous construction processor with the deterministic reconstruction baseline (homo. c. + det. r.), despite the fact they were not trained together.
@@ 3.1 Results
Table 1: In- and out- of distribution performance for different models. Best results are boldfaced, but we separate models using an oracle in their execution.
$n$ $C$ no-hint baseline reg. c. + reg. r. homo. c. + reg. r. homo. c. + det. r.
micro-F1 exact micro-F1 exact micro-F1 exact micro-F1 exact
$16$ $16$ $0.889_{\scriptscriptstyle\pm 0.017}$ $0.385_{\scriptscriptstyle\pm 0.101}$ $\mathbf{0.989}_{\scriptscriptstyle\pm 0.004}$ $\mathbf{0.906}_{\scriptscriptstyle\pm 0.031}$ $0.985_{\scriptscriptstyle\pm 0.007}$ $0.901_{\scriptscriptstyle\pm 0.033}$ $0.975_{\scriptscriptstyle\pm 0.004}$ $0.891_{\scriptscriptstyle\pm 0.013}$
$16$ $64$ $\mathbf{0.943}_{\scriptscriptstyle\pm 0.000}$ $0.250_{\scriptscriptstyle\pm 0.000}$ $0.785_{\scriptscriptstyle\pm 0.063}$ $0.438_{\scriptscriptstyle\pm 0.095}$ $0.917_{\scriptscriptstyle\pm 0.017}$ $\mathbf{0.495}_{\scriptscriptstyle\pm 0.115}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.957}}_{\scriptscriptstyle\pm 0.012}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.578}}_{\scriptscriptstyle\pm 0.109}$
$32$ $32$ $656_{\scriptscriptstyle\pm 0.122} 0.0$$000_{\scriptscriptstyle\pm 0.000}$ $0.854_{\scriptscriptstyle\pm 0.015}$ $0.177_{\scriptscriptstyle\pm 0.039}$ $\mathbf{0.925}_{\scriptscriptstyle\pm 0.040}$ $\mathbf{0.510}_{\scriptscriptstyle\pm 0.173}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.933}}_{\scriptscriptstyle\pm 0.034}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.521}}_{\scriptscriptstyle\pm 0.128}$
$64$ $16$ $835_{\scriptscriptstyle\pm 0.030} 0.0$$031_{\scriptscriptstyle\pm 0.031}$ $0.824_{\scriptscriptstyle\pm 0.055}$ $0.135_{\scriptscriptstyle\pm 0.127}$ $\mathbf{0.939}_{\scriptscriptstyle\pm 0.011}$ $\mathbf{0.615}_{\scriptscriptstyle\pm 0.110}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.949}}_{\scriptscriptstyle\pm 0.015}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.635}}_{\scriptscriptstyle\pm 0.102}$
$64$ $64$ $543_{\scriptscriptstyle\pm 0.000} 0.0$$000_{\scriptscriptstyle\pm 0.000}$ $448_{\scriptscriptstyle\pm 0.084} 0.0$$000_{\scriptscriptstyle\pm 0.000}$ $\mathbf{0.668}_{\scriptscriptstyle\pm 0.087}$ $0.000_{\scriptscriptstyle\pm 0.000}$ $\mathbf{0.770}$±0.115 $0.000_{\scriptscriptstyle\pm 0.000}$
@@
The second result we want to highlight is the increase in performance when using a homogeneous NAR constructor. In the larger OOD regimes (homo. c. + reg. r.) significantly outperforms the baseline F1 scores and achieves best exact-match accuracy from the fully-neural models. According to our observations, even if the input value scales remain fixed across test distributions, homogenity could be one of the requirements when modelling pseudo-polynomial DP problems, since the magnitude of the solution increases with size (Appendix F).
@@
In spite of the solid results, there are limitations remaining to be addressed: 1) Oracle instability – we were unable to train (homo. c. + det. r.), due to exploding/vanishing gradients. We hypothesise this is due to the algorithm chaining multiplications of probabilities and/or operating in scalar space (Appendix G). 2) Separated trainings – currently, the training of the construction model is completely decoupled from the training of the reconstruction one. This is not common for NAR models and is a topic of our future interests. Looking ahead, we also plan to investigate how the observations made in this work, such as the use of edge length encoding, could be applied to problems from CLRS-30, shedding light on their potential beyond the current setting.
@@ Appendix A Related work
Neural algorithmic reasoning aligns learned computation with classical algorithms, with the aim of obtaining trained neural models that can execute the target algorithm in the out-of-distribution (i.e. larger input instances) regime. Intermediate-step supervision [11, 2], also utilised here, has been the initial approach to achieve such alignment. However, this alone is not sufficient and architectural modifications were necessary: Early analyses studied what neural networks can reason about and how they extrapolate [12, 13] – on an informal level their findings state that neural networks extrapolate well if the algorithm can be separated into subfunctions and each of them is easily learnable by a corresponding neural submodule. This, combined with the rise of standardised NAR benchmarks (CLRS-30 [2] and SALSA-CLRS [14]) gave rise to many subsequent works, proposing architectural alignment strategies such as homogeneity [8], persistent message passing [15], relational/positional attention [16, 17], and recurrent aggregators [18]. Related efforts explore structured memory for recursion and data structures [19, 20], while others investigate reducing or re-structuring hints [21, 10, 22, 23, 24].
@@
Graph neural networks [25], the underlying architecture behind the NAR processors, are theoretically aligned to dynamic programming algorithms [9]. However, when NAR is benchmarked, most studies address polynomial-time DPs, whereas the 011 Knapsack algorithm is pseudo-polynomial (weakly NP-hard) and was deliberately omitted from CLRS-30. Outside of the field of NAR, there exist DP-inspired neural networks [26] that provably solve Knapsack, but their weights are hardcoded, rather than learned from data, as is the case in our work. Being weakly NP-hard, classical neural combinatorial optimisation (NCO) approaches could also be applied, but usually they rely on reinforcement learning [27, 28, 29, 30], which we aimed to avoid. From the NCO approaches, our work is most conceptually similar to those that supervise on input-output mapping [31, 32, 33]. Still, we differ by explicitly mirroring the pseudo-polynomial DP algorithm in a two-phase construction–reconstruction pipeline, supervising intermediate states and forcing table-guided backtracking, which improves out-of-distribution generalization.

### v2

@@ 1 Introduction
Knapsack is NP-hard. [4] CLRS-30 omits this class of problems, due to the impossibility of generating sufficiently many samples for them (unless NP=co-NP) [5]. However, for Knapsack, there exists a dynamic programming solution whose runtime depends on the numeric value of the capacity rather than its bit-length. This makes Knapsack a pseudo-polynomial problem [^1] Informal term meaning the problem admits a pseudo-polynomial algorithm; formally, it is weakly NP-hard., lifting the implications of Yehuda et al. [5], provided we do not consider extreme capacities. Despite its theoretical and practical significance, the problem has not been explored within the NAR framework. Thus, we develop a neural algorithmic reasoner for Knapsack following a two-phase pipeline: dynamic programming table construction and solution reconstruction, as shown in Figure 1. This approach is transferable to other pseudo-polynomial problems, as further discussed in Appendix B.
@@ 2.2 Pseudo-Polynomial Modeling
where $\delta$ is the Kronecker delta function and $M=10$ is the cutoff. Such categorical encoding gives the model the ability to choose the correct past states that influence the current one. Edge length encoding is illustrated in Figure 2. Figure 3 shows the impact of EL on an in-distribution problem instance in comparison to standard random node positional encoding. For more details, see Appendix F.
We additionally observe that the CLRS-30’s processors struggle to handle the Knapsack DP scalar values larger than those seen in training. We propose to mitigate this issue by adopting a homogeneous processor [8], which enforces the model to be invariant to the scale of the item values. This substantially improves the value generalization performance of the model on instances larger than those seen in training. Appendix G provides more results and a discussion.
@@
where $\bm{p}$} and $\bm{w}$} are the vectors of positional encodings and weights (capacity nodes are padded with 0), $0\leq c,c^{\prime}\leq C$ and $C+1\leq i\leq C+n$ and ${\mathbb{I}}_{\mathcal{G}}=\emptyset$. To model the reconstruction process, the hints provide information on the item currently under consideration, the total remaining capacity, and the set of items selected up to that point. For each item, the model outputs the probability that it belongs to the solution. Due to the space constraints, details can be found in Appendix E.
To reconstruct the solution from the $\mathrm{decision}$ table, it is sufficient to have information about the item weights. During reconstruction, as shown in Figure 4, having item values leads to worse generalization, as the model is trying to find a shortcut to the solution, without actually learning to navigate the $\mathrm{decision}$ table. We hypothesise that this is the reason why the joint model approach did not work. We also note that having two alternating steps, one for decision and one for moving the pointer works better than combining the two. Further details are provided in Figure 8, Appendix E.
@@
An alternative would be to perform reconstruction based on the $\mathrm{decision}$ probability table, as it allows for end-to-end differentiability. In Appendix H, we describe in more detail such a differentiable deterministic reconstruction. Sadly, our initial experiments in training an NAR construction, coupled with deterministic reconstruction, were unstable. Nevertheless, in the evaluation, we include as a baseline a model that, during the inference phase, performs deterministic reconstruction based on the $\mathrm{decision}$ table generated by NAR construction model.
@@ 3 Evaluation
Our experiments are implemented on top of CLRS-30 [7] (3 seeds for std). Hyperparameters (learning rate, optimiser, etc.) are inherited, including the use of Triplet-GMPNN processor [9] as well as the standard node positional encoding. Differently from CLRS-30, we use the model from the final epoch for testing, as it was recommended in [10]. Furthermore, during the construction training phase, we perform 30,000 epochs, as we observed that the model does not stabilize within the first 10,000 epochs.
The model is trained on samples where $n\leq 16$ and $C\leq 16$ (in-distribution). For out-of-distribution experiments, we consider different values of $n$ and $C$ (up to $n=64$ and $C=64$). Each configuration consists of 64 samples. Item weights are uniformly sampled from $\{1,2,\ldots,8\}$, and item values are uniformly sampled from the interval $[0,1]$.
Following CLRS-30, during testing, we report the micro-F1 score. We also add exact-match accuracy – the fraction of samples whose discretized solution, obtained by greedily selecting the highest-probability item until capacity is reached, exactly matches the ground truth.
@@
Alongside the models with regular processors (reg. c. + reg. r.), and a homogeneous processor during construction (homo. c. + reg. r.), we evaluate a no-hint baseline, which directly predicts the optimal subset from the input. It is implemented using the no-hint mode in CLRS-30 with $2n$ message-passing steps, and includes scalar features for item weights and capacity. We additionally zero-shot evaluate the homogeneous construction processor with the deterministic reconstruction baseline (homo. c. + det. r.), despite the fact they were not trained together. We also compare our models to Tropical Attention [11], a recent direct-prediction model for combinatorial optimization problems based on tropical projective geometry and polyhedral geometric inductive bias. To make the comparison, we adjusted the sampler and the training and test sizes to match ours, see Appendix C. For more information on this and other relevant works, see Appendix A.
@@ 3.1 Results
Table 1: In- and out- of distribution performance for different models. Best results are boldfaced, but we separate models using an oracle in their execution. Best viewed on screen.
$n$ $C$ no-hint baseline reg. c. + reg. r. homo. c. + reg. r. homo. c. + det. r. Tropical Attention [11]
micro-F1 exact micro-F1 exact micro-F1 exact micro-F1 exact micro-F1 exact
$16$ $16$ $0.889_{\scriptscriptstyle\pm 0.017}$ $0.385_{\scriptscriptstyle\pm 0.101}$ $\mathbf{0.989}_{\scriptscriptstyle\pm 0.004}$ $\mathbf{0.906}_{\scriptscriptstyle\pm 0.031}$ $0.985_{\scriptscriptstyle\pm 0.007}$ $0.901_{\scriptscriptstyle\pm 0.033}$ $0.975_{\scriptscriptstyle\pm 0.004}$ $0.891_{\scriptscriptstyle\pm 0.013}$ $0.909_{\scriptscriptstyle\pm 0.013}$ $0.381_{\scriptscriptstyle\pm 0.005}$
$16$ $64$ $\mathbf{0.943}_{\scriptscriptstyle\pm 0.000}$ $0.250_{\scriptscriptstyle\pm 0.000}$ $0.785_{\scriptscriptstyle\pm 0.063}$ $0.438_{\scriptscriptstyle\pm 0.095}$ $0.917_{\scriptscriptstyle\pm 0.017}$ $\mathbf{0.495}_{\scriptscriptstyle\pm 0.115}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.957}}_{\scriptscriptstyle\pm 0.012}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.578}}_{\scriptscriptstyle\pm 0.109}$ $0.892_{\scriptscriptstyle\pm 0.011}$ $0.493_{\scriptscriptstyle\pm 0.027}$
$32$ $32$ $656_{\scriptscriptstyle\pm 0.122} 0.0$$000_{\scriptscriptstyle\pm 0.000}$ $0.854_{\scriptscriptstyle\pm 0.015}$ $0.177_{\scriptscriptstyle\pm 0.039}$ $\mathbf{0.925}_{\scriptscriptstyle\pm 0.040}$ $\mathbf{0.510}_{\scriptscriptstyle\pm 0.173}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.933}}_{\scriptscriptstyle\pm 0.034}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.521}}_{\scriptscriptstyle\pm 0.128}$ $0.923_{\scriptscriptstyle\pm 0.012}$ $0.186_{\scriptscriptstyle\pm 0.078}$
$64$ $16$ $835_{\scriptscriptstyle\pm 0.030} 0.0$$031_{\scriptscriptstyle\pm 0.031}$ $0.824_{\scriptscriptstyle\pm 0.055}$ $0.135_{\scriptscriptstyle\pm 0.127}$ $\mathbf{0.939}_{\scriptscriptstyle\pm 0.011}$ $\mathbf{0.615}_{\scriptscriptstyle\pm 0.110}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.949}}_{\scriptscriptstyle\pm 0.015}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.635}}_{\scriptscriptstyle\pm 0.102}$ $822_{\scriptscriptstyle\pm 0.024} 0.0$$043_{\scriptscriptstyle\pm 0.047}$
$64$ $64$ $543_{\scriptscriptstyle\pm 0.000} 0.0$$000_{\scriptscriptstyle\pm 0.000}$ $448_{\scriptscriptstyle\pm 0.084} 0.0$$000_{\scriptscriptstyle\pm 0.000}$ $\mathbf{0.668}_{\scriptscriptstyle\pm 0.087}$ $0.000_{\scriptscriptstyle\pm 0.000}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.770}}_{\scriptscriptstyle\pm 0.115}$ $0.000_{\scriptscriptstyle\pm 0.000}$ $374_{\scriptscriptstyle\pm 0.000} 0.0$$000_{\scriptscriptstyle\pm 0.000}$
@@
The second result we want to highlight is the increase in performance when using a homogeneous NAR constructor. In the larger OOD regimes (homo. c. + reg. r.) significantly outperforms the baseline F1 scores and achieves best exact-match accuracy from the fully-neural models. According to our observations, even if the input value scales remain fixed across test distributions, homogeneity could be one of the requirements when modelling pseudo-polynomial DP problems, since the magnitude of the solution increases with size (Appendix G).
@@
Our best fully neural model (homo. c. + reg. r.) achieves better performance than Tropical Attention in all OOD settings, but it is important to interpret these results in the context of the differences between approaches. We note that Tropical Attention’s primary aim, like ours, is not to compete with traditional or neural Knapsack solvers, but rather to demonstrate the effectiveness of their proposed architecture. We further hypothesize that combining Tropical Attention’s insights with ours – which are largely orthogonal – could yield even stronger performance.
3.2 Properties of Generated Scalar DP Tables
In Table 5, Appendix G, we report qualities of the generated $\mathrm{decision}$ tables, but – aside from a few visual examples – we do not quantify the quality of the predicted scalar DP tables. Because mean squared error is not very interpretable for this purpose, we perform an analysis based on three statistics that correspond to the fundamental properties of the Knapsack DP table. To avoid issues with numerical precision of predicted values, we use $\varepsilon=0.01$, which is sufficiently precise for gaining insights into the constructed DP tables, given that item values are sampled from $[0,1]$. These statistics measure the proportion of DP table cells for which the specified constraints hold.
Table 2: Proportions of DP table entries satisfying three fundamental properties under $\varepsilon{=}0.01$: (1) item-wise monotonicity: $\mathrm{dp}_{i,c}\geq\mathrm{dp}_{i-1,c}-\varepsilon$; (2) capacity-wise monotonicity: $\mathrm{dp}_{i,c}\geq\mathrm{dp}_{i,c-1}-\varepsilon$; (3) optimal substructure: $\mathrm{dp}_{i,c}\geq\max(\mathrm{dp}_{i-1,c},\mathrm{dp}_{i-1,c-w_{i}}+v_{i})-\varepsilon$.
$n$ $C$ item-wise mon. capacity-wise mon. optimal substructure
reg. c. homo. c. reg. c. homo. c. reg. c. homo. c.
$16$ $16$ $0.893_{\scriptscriptstyle\pm 0.032}$ $\mathbf{0.974}_{\scriptscriptstyle\pm 0.009}$ $\mathbf{0.951}_{\scriptscriptstyle\pm 0.027}$ $0.945_{\scriptscriptstyle\pm 0.036}$ $0.700_{\scriptscriptstyle\pm 0.057}$ $\mathbf{0.948}_{\scriptscriptstyle\pm 0.015}$
$16$ $64$ $0.915_{\scriptscriptstyle\pm 0.049}$ $\mathbf{0.988}_{\scriptscriptstyle\pm 0.004}$ $\mathbf{0.892}_{\scriptscriptstyle\pm 0.035}$ $0.799_{\scriptscriptstyle\pm 0.062}$ $0.377_{\scriptscriptstyle\pm 0.070}$ $\mathbf{0.677}_{\scriptscriptstyle\pm 0.043}$
$64$ $16$ $0.711_{\scriptscriptstyle\pm 0.014}$ $\mathbf{0.978}_{\scriptscriptstyle\pm 0.003}$ $0.975_{\scriptscriptstyle\pm 0.010}$ $\mathbf{0.981}_{\scriptscriptstyle\pm 0.014}$ $0.548_{\scriptscriptstyle\pm 0.031}$ $\mathbf{0.967}_{\scriptscriptstyle\pm 0.006}$
$32$ $32$ $0.773_{\scriptscriptstyle\pm 0.028}$ $\mathbf{0.983}_{\scriptscriptstyle\pm 0.004}$ $\mathbf{0.967}_{\scriptscriptstyle\pm 0.009}$ $0.967_{\scriptscriptstyle\pm 0.018}$ $0.447_{\scriptscriptstyle\pm 0.060}$ $\mathbf{0.901}_{\scriptscriptstyle\pm 0.028}$
$64$ $64$ $0.636_{\scriptscriptstyle\pm 0.022}$ $\mathbf{0.981}_{\scriptscriptstyle\pm 0.005}$ $\mathbf{0.860}_{\scriptscriptstyle\pm 0.031}$ $0.841_{\scriptscriptstyle\pm 0.064}$ $0.241_{\scriptscriptstyle\pm 0.032}$ $\mathbf{0.715}_{\scriptscriptstyle\pm 0.036}$
The results in Table 2 show that the homogeneous processor model produces significantly better tables in terms of item-wise monotonicity and optimal substructure property, while capacity-wise monotonicity results are comparable. The results suggest that capacity generalization presents a greater challenge than item count generalization when predicting DP scalars. While the optimal substructure property was violated 30% of the time for our largest OOD experiment, the 46% improvement was sufficient to improve our downstream predictions. Note that there exist other alternative approaches [12], that provably generalise to any input size, but those mimic the algorithms too closely, trading the flexibility of neural networks, which is one of the reasons NAR exists.
@@ 3.2 Properties of Generated Scalar DP Tables
In spite of the solid results, there are limitations remaining to be addressed: 1) Oracle instability – we were unable to train (homo. c. + det. r.), due to exploding/vanishing gradients. We hypothesise this is due to the algorithm chaining multiplications of probabilities and/or operating in scalar space (Appendix H). 2) Separated trainings – currently, the training of the construction model is completely decoupled from the training of the reconstruction one. This is not common for NAR models and is a topic of our future interests. Looking ahead, we also plan to investigate how the observations made in this work, such as the use of edge length encoding, could be applied to problems from CLRS-30, shedding light on their potential beyond the current setting.
@@ Appendix A Related Work
Neural algorithmic reasoning aligns learned computation with classical algorithms, with the aim of obtaining trained neural models that can execute the target algorithm in the out-of-distribution regime. Intermediate-step supervision [13, 2], also utilised here, has been the initial approach to achieve such alignment. However, this alone is not sufficient and architectural modifications were necessary. Early analyses studied what neural networks can reason about and how they extrapolate [14, 15] – on an informal level their findings state that neural networks extrapolate well if the algorithm can be separated into subfunctions and each of them is easily learnable by a corresponding neural submodule. This, combined with the rise of standardised NAR benchmarks (CLRS-30 [2] and SALSA-CLRS [16]) gave rise to many subsequent works, proposing architectural alignment strategies such as homogeneity [8], persistent message passing [17], relational/positional attention [18, 19], and recurrent aggregators [20]. Related efforts explore structured memory for recursion and data structures [21, 22], while others investigate reducing or re-structuring hints [23, 10, 24, 25, 26].
@@
Graph neural networks [27], the underlying architecture behind the NAR processors, are theoretically aligned to dynamic programming algorithms [9]. However, when NAR is benchmarked, most studies address polynomial-time DPs, and the 011 Knapsack algorithm, being pseudo-polynomial (weakly NP-hard), was deliberately omitted from CLRS-30. While some works have addressed NP-hard problems [28, 29], none have focused specifically on pseudo-polynomial problems. Outside of the field of NAR, there exist DP-inspired neural networks [30] that provably solve Knapsack, but their weights are hardcoded, rather than learned from data, as is the case in our work. Due to the nature of our problem, classical neural combinatorial optimisation (NCO) approaches could also be applied. Some possible approaches are Bello et al. [31], Kwon et al. [32], Afshar et al. [33], Zhang et al. [34], but these are fundamentally different from our approach. Moreover, these approaches represent item values and weights as real numbers and use fixed capacity. Additionally, they do not consider out-of-distribution settings at all – the instance size is fixed and ranges up to several hundred items. It is worth noting that the idea of combining NAR and RL is very recent [35], and there are no approaches that would tackle pseudo-polynomial algorithms in this manner.
From the NCO approaches, our work is most conceptually similar to supervised methods that solve Knapsack by learning the input-output mapping directly, without following the algorithm’s intermediate steps [36, 37, 11]. One supervised approach [36], which uses the same sampler as Kwon et al. [32], does consider OOD, but trains on sizes of $100$ and tests on $200$, $500$, and $1000$, which is not comparable to our setting. Another supervised learning approach for solving Knapsack [37] does not have available source code and reproducibility is not at the required level. The only work we found that is comparable to ours in terms of the training and testing scales is Tropical Attention [11]. Tropical Attention [11] introduces a solution based on tropical projective geometry and polyhedral geometric inductive bias. The Tropical Transformers enhance OOD performance in both length and value generalization on several combinatorial optimization problems. In the Knapsack experiments, the model is trained for $n=8$ and tested for length generalization to $n=64$. Capacity generalization is not performed – capacity is integer from $10$ to $20$, in both cases. The model has much shorter training time, primarily due to the lightweight implementation and the fact that it does not use hints, but rather performs supervised direct-prediction of outputs based on inputs. The source code for Tropical Attention is available, so we adapted the sampler and training and test sizes to compare it with our approach.